OpenStudy (anonymous):

A vector v has initial point (22, 25) and terminal point ( -11, 13). Find its magnitude and direction

OpenStudy (anonymous):

v = <-33,12> ||v|| = sqrt [ 33^2 + 12^2 ] which is sqrt 265 yes... but is that the answer??

OpenStudy (anonymous):

@surjithayer ^

OpenStudy (anonymous):

what about the direction?

OpenStudy (anonymous):

vector v= (-11-22)i+(13-25)j=-33i-12j \[magnitude=\left| v \right|=\left| \sqrt{\left( -33 \right)^{2}+\left( -12 \right)^{2}} \right|\] \[=\sqrt{1089+144}=\sqrt{1233}=3\sqrt{137}\] \[unit~ normal=\frac{ v }{ \left| v \right| }=\frac{ -33i-12j }{ 3\sqrt{137} }=\frac{-11i-4j }{ \sqrt{137} }\]

OpenStudy (anonymous):

thats the unit vector in direction of v.... but its not asking for a unit vector.. just a regular vector @surjithayer

OpenStudy (anonymous):

correction. it is unit vector not unit normal.

OpenStudy (anonymous):

so how do I find the direction of the vector without reducing its size down to 1 unit?

OpenStudy (anonymous):

v is regular vector.

OpenStudy (anonymous):

so... to answer the question.. I just give the magnitude which is 3 sqrt(137) and v = <-33, 12> for direction?

OpenStudy (anonymous):

the direction is the part of the answer that confuses me... This is a practice exam for studying purposes, but i really want to understand this stuff so i can ace the exam tomorow

OpenStudy (anonymous):

v=<-33,-12> for direction we generally give direction ratios or direction cosines. here d.r's are <-33,-12> d.c's are <-11/sqrt137,-4/sqrt137>

OpenStudy (anonymous):

oh okay.. so that explains it then. Thankyou very much sir... I really appreciate it

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so its <a2/||v|| , b2/ ||v|| > right ??